OpenStudy (anonymous):

A boat traveled 100 miles against the current in 5 hours. The return trip downstream took 4 hours. What is the speed of the boat in still water? What is the rate of the current?

OpenStudy (anonymous):

Let, speed of boat x miles / hr, speed of current = y Hence, upstream, net speed (x-y) mi / hr, hence, 100/ (x-y) = 5 => x - y= 20, downstream, net speed (x+y) mi /hr, hence, 100/(x+y) = 4 => x+y = 25 By elimination, x = 22.5 mi/hr, y = 2.5 mi/ hr

OpenStudy (anonymous):

80 miles

OpenStudy (anonymous):

i think you divide 100/5 = 20 20 * 4 = 80

OpenStudy (anonymous):

I am sorry, but why'd you multiply by 4 again